OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

myininaya (myininaya):

so we are to find f(g(-1)) so we notice the inside function is g and so we want to find g(-1) first

myininaya (myininaya):

To find g(-1) you will need to use g(x)=x^2-7x-9

myininaya (myininaya):

notice we want the input to be -1 so we replace all the x's with 1's in g(x)=x^2-7x-9 can you do that?

myininaya (myininaya):

-1's*

OpenStudy (anonymous):

=1^2-7x-9

myininaya (myininaya):

you need to replace all the x's in g(x)=x^2-7x-9 with -1

myininaya (myininaya):

not just some

myininaya (myininaya):

so you should have \[g(-1)=(-1)^2-7(-1)-9\] can you simplify that?

OpenStudy (anonymous):

-2-(-7)-9

myininaya (myininaya):

(-1)^2=(-1)(-1) this is not the same as -1-1

myininaya (myininaya):

or however your reading it it is not the same as (-1)(2) (-1)^2 means (-1)(-1)

OpenStudy (jdoe0001):

\(\bf f(x) = x - 2 \qquad {\color{brown}{ g(x)}} = x^2 - 7x - 9 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )={\color{brown}{ g(x)}}-2\implies f(\quad {\color{brown}{ g(x)}}\quad )=({\color{brown}{ x^2 - 7x - 9}})-2 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ -1}})}}\quad )=({\color{brown}{ ({\color{blue}{ -1}})^2 - 7({\color{blue}{ -1}}) - 9}})-2\)